OpenStudy (anonymous):

@phi find the sum 2 root 6, root 24

OpenStudy (phi):

\[ 2\sqrt{6} + \sqrt{24} \] the idea is the sqr(24) might be simplified. I would factor the 24 and look for pairs of numbers to "pull out" of the square root.

OpenStudy (anonymous):

4 root of 3

OpenStudy (phi):

how ?

OpenStudy (akashdeepdeb):

2 rt 6 is actually same as rt 24 So? When you add them you get their double! :) Ans: 4 rt 6! :D

OpenStudy (anonymous):

nvm could I have a answer choice of being 2 root6, 4 root 3!

OpenStudy (phi):

are you saying you can't notice that 24 is 2*12, or 2*2*6 (we could go further) but sqr(2*2*6) has a pair of 2's you can pull out the pair, and write 2 in front: 2sqr(6) now you can combine like terms

OpenStudy (anonymous):

13 root 6?

OpenStudy (phi):

\[ 2\sqrt{6} + \sqrt{24} \\ 2\sqrt{6} + \sqrt{2\cdot 2\cdot 6} \\2\sqrt{6} + 2\sqrt{6}\]

OpenStudy (phi):

count how many sqr(6)'s you have : 2+2 is ?

OpenStudy (phi):

you do know \[ 2 \sqrt{6} \text{ means } \sqrt{6}+\sqrt{6} \] you have 2 sqr(6)'s add 2 more how many do you have all together ?

OpenStudy (anonymous):

4 root 6

OpenStudy (anonymous):

@mary.rojas

OpenStudy (phi):

yes

OpenStudy (mary.rojas):

What are you stuck on? all? some?

OpenStudy (anonymous):

I don't understand the problem at all

OpenStudy (phi):

do you understand that \[ 2\sqrt{6}+2\sqrt{6} = 4\sqrt{6}\] ?

OpenStudy (anonymous):

yes, just a little bit, sorry for taking time im a visual learner and its hard to read it and understand it

OpenStudy (phi):

you should think about this general idea: 2 of "something" plus 2 more of "something" is 4 "something" that works for something being cows, horses, (x+3), \( \sqrt{6} \) and so on

OpenStudy (mary.rojas):